OpenStudy (anonymous):

For the substitution method do you put the variable in front? ex: 4r+s=19 -4r -4r s=?

OpenStudy (the_fizicx99):

What do you mean substitution method, you're subtracting 4r from both sides, why?

OpenStudy (anonymous):

I have to use the substitiution method to answer this question: \[5r+2s=23 \] \[4r+s=19\]

OpenStudy (anonymous):

\[4r+s=19\] \[4r=19-s\] \[4r-19=-s\] \[\frac{ 4r-19 }{ -1 }=\frac{ -s }{ -1 }\] \[s=-4r+19\] Plug in for s in first equation. \[5r+(-4r+19)=23\] \[5r+-4r+19=23\] \[r+19=23\] Subtract 19 from 23 \[r=4\]